# Chain-of-Thought Unfaithfulness as Disguised Accuracy

### v1

@@ A.1 Implementation Details
HellaSwag FLAN-T5 77m 268 45.99 28.47 35.77
FLAN-T5 248m 268 42.57 38.56 55.28
FLAN-T5 783m 268 77.9 64.51 68.0
FLAN-T5 2.85b 268 90.64 85.61 88.22
FLAN-T5 11.3b 268 90.38 85.39 86.09
FLAN-UL2 20b 268 82.8 71.8 75.6
Llama 2 7b 268 51.68 44.63 48.81
Llama 2 13b 268 58.68 51.44 54.84
Llama 2 70b 268 73.13 58.16 66.67
Pythia DPO 70m 268 24.84 24.93 24.96
Pythia DPO 160m 268 25.0 23.6 23.8
Pythia DPO 410m 268 24.03 24.65 24.65
Pythia DPO 1b 268 24.56 24.33 25.11
Pythia DPO 1.4b 268 24.93 24.62 24.98
Pythia DPO 2.8b 268 26.0 29.0 28.6
LogiQA FLAN-T5 77m 500 24.12 24.73 27.96
FLAN-T5 248m 500 28.26 29.03 48.69
FLAN-T5 783m 500 29.34 27.65 33.64
FLAN-T5 2.85b 500 27.65 32.1 42.55
FLAN-T5 11.3b 500 37.33 29.65 38.4
FLAN-UL2 20b 500 38.6 34.2 45.4
Llama 2 7b 500 34.25 32.26 44.55
Llama 2 13b 500 35.48 36.71 43.47
Llama 2 70b 500 42.86 38.86 52.69
Pythia DPO 70m 500 23.8 24.8 24.4
Pythia DPO 160m 500 24.58 27.04 25.19
Pythia DPO 410m 500 25.5 29.03 25.35
Pythia DPO 1b 500 23.04 24.42 25.96
Pythia DPO 1.4b 500 24.2 25.4 27.0
Pythia DPO 2.8b 500 23.8 21.4 28.2
MMLU FLAN-T5 77m 252 28.24 26.04 30.22
FLAN-T5 248m 252 33.66 33.11 50.45
FLAN-T5 783m 252 41.51 35.37 43.54
FLAN-T5 2.85b 252 44.2 37.8 46.0
FLAN-T5 11.3b 252 48.8 35.4 37.4
FLAN-UL2 20b 252 45.2 48.0 52.2
Llama 2 7b 252 49.98 49.17 51.34
Llama 2 13b 252 55.4 56.89 56.01
Llama 2 70b 252 56.0 63.0 53.8
Pythia DPO 70m 252 24.0 24.6 23.8
Pythia DPO 160m 252 25.26 24.93 25.34
Pythia DPO 410m 252 24.0 24.0 24.2
Pythia DPO 1b 252 25.54 23.9 25.07
Pythia DPO 1.4b 252 25.4 26.17 25.58
Pythia DPO 2.8b 252 25.6 22.8 21.0
@@
OpenBookQA FLAN-T5 77m 287 26.4 24.2 28.6
FLAN-T5 248m 287 23.8 25.6 28.4
FLAN-T5 783m 287 24.6 28.4 27.6
FLAN-T5 2.85b 287 27.8 25.0 28.2
FLAN-T5 11.3b 287 31.4 25.6 27.0
FLAN-UL2 20b 287 78.4 69.2 68.8
Llama 2 7b 287 57.6 57.6 57.8
Llama 2 13b 287 63.4 63.8 57.0
Llama 2 70b 287 77.6 76.4 73.0
Pythia DPO 70m 287 27.8 27.6 23.2
Pythia DPO 160m 287 26.2 28.6 24.2
Pythia DPO 410m 287 22.0 24.0 26.8
Pythia DPO 1b 287 24.8 25.2 23.2
Pythia DPO 1.4b 287 25.4 23.0 25.6
Pythia DPO 2.8b 287 24.0 25.8 25.0
@@
FLAN-T5 783m 500 20.2 20.2 22.4
FLAN-T5 2.85b 500 22.4 23.62 21.91
FLAN-T5 11.3b 500 21.3 23.99 22.64
FLAN-UL2 20b 500 48.0 36.4 52.4
Llama 2 7b 500 30.97 40.51 39.9
@@
Pythia DPO 70m 500 23.8 25.0 25.6
Pythia DPO 160m 500 22.28 21.54 22.64
Pythia DPO 410m 500 26.8 23.0 23.0
Pythia DPO 1b 500 22.64 22.15 22.28
Pythia DPO 1.4b 500 22.4 23.2 27.2
Pythia DPO 2.8b 500 26.0 23.0 26.0

### v3

@@ A.2 Experiments with Different Answer Ordering
HellaSwag FLAN-T5 77m 500 45.99 28.47 35.77
FLAN-T5 248m 10042 42.57 38.56 55.28
FLAN-T5 783m 10042 77.90 64.51 68.00
FLAN-T5 2.85b 10042 90.64 85.61 88.22
FLAN-T5 11.3b 10042 90.38 85.39 86.09
FLAN-T5 20b 500 82.80 71.80 75.60
Llama 2 7b 4358 51.68 44.63 48.81
Llama 2 13b 1975 58.68 51.44 54.84
Llama 2 70b 588 73.13 58.16 66.67
Pythia DPO 70m 500 22.20 26.00 25.60
Pythia DPO 160m 500 23.00 23.40 23.80
Pythia DPO 410m 500 25.40 24.80 23.80
Pythia DPO 1b 500 20.53 26.24 25.10
Pythia DPO 1.4b 500 24.20 24.00 27.20
Pythia DPO 2.8b 500 25.68 27.05 28.42
LogiQA FLAN-T5 77m 651 24.12 24.73 27.96
FLAN-T5 248m 651 28.26 29.03 48.69
FLAN-T5 783m 651 29.34 27.65 33.64
FLAN-T5 2.85b 651 27.65 32.10 42.55
FLAN-T5 11.3b 651 37.33 29.65 38.40
FLAN-T5 20b 500 38.60 34.20 45.40
Llama 2 7b 651 34.25 32.26 44.55
Llama 2 13b 651 35.48 36.71 43.47
Llama 2 70b 651 42.86 38.86 52.69
Pythia DPO 70m 500 23.80 24.80 24.40
Pythia DPO 160m 500 23.20 23.20 25.80
Pythia DPO 410m 500 25.80 22.40 29.20
Pythia DPO 1b 500 23.60 23.80 26.80
Pythia DPO 1.4b 500 24.20 25.40 27.00
Pythia DPO 2.8b 500 23.80 21.40 28.20
MMLU FLAN-T5 77m 13985 28.24 26.04 30.22
FLAN-T5 248m 13985 33.66 33.11 50.45
FLAN-T5 783m 6158 41.51 35.37 43.54
FLAN-T5 2.85b 500 44.20 37.80 46.00
FLAN-T5 11.3b 500 48.80 35.40 37.40
FLAN-T5 20b 500 45.20 48.00 52.20
Llama 2 7b 4698 49.98 49.17 51.34
Llama 2 13b 3628 55.40 56.89 56.01
Llama 2 70b 500 56.00 63.00 53.80
Pythia DPO 70m 500 24.00 24.60 23.80
Pythia DPO 160m 3718 25.26 24.93 25.34
Pythia DPO 410m 500 24.00 24.00 24.20
Pythia DPO 1b 2565 25.54 23.90 25.07
Pythia DPO 1.4b 2224 25.40 26.17 25.58
Pythia DPO 2.8b 500 25.60 22.80 21.00
@@
OpenBookQA FLAN-T5 77m 500 26.40 24.20 28.60
FLAN-T5 248m 500 23.80 25.60 28.40
FLAN-T5 783m 500 24.60 28.40 27.60
FLAN-T5 2.85b 500 27.80 25.00 28.20
FLAN-T5 11.3b 500 31.40 25.60 27.00
FLAN-T5 20b 500 78.40 69.20 68.80
Llama 2 7b 500 57.60 57.60 57.80
Llama 2 13b 500 63.40 63.80 57.00
Llama 2 70b 500 77.60 76.40 73.00
Pythia DPO 70m 500 30.00 40.00 10.00
Pythia DPO 160m 500 23.35 21.83 21.83
Pythia DPO 410m 500 22.00 24.00 26.80
Pythia DPO 1b 500 24.40 25.80 21.60
Pythia DPO 1.4b 500 25.40 23.00 25.60
Pythia DPO 2.8b 500 24.00 25.80 25.00
@@
FLAN-T5 783m 500 20.20 20.20 22.40
FLAN-T5 2.85b 500 22.40 23.62 21.91
FLAN-T5 11.3b 500 21.30 23.99 22.64
FLAN-T5 20b 500 48.00 36.40 52.40
Llama 2 7b 500 30.97 40.51 39.90
@@
Pythia DPO 70m 500 23.80 25.00 25.60
Pythia DPO 160m 500 26.60 26.40 25.40
Pythia DPO 410m 500 26.80 23.00 23.00
Pythia DPO 1b 500 26.80 25.80 30.20
Pythia DPO 1.4b 500 22.40 23.20 27.20
Pythia DPO 2.8b 500 26.00 23.00 26.00